OpenStudy (anonymous):

What is a possible first step to eliminate a variable in the following system? 2x - 3y = -1 3x - 4y = -3

OpenStudy (anonymous):

a) Multiply the first equation by -3 and the second equation by 2. b) Multiply the first equation by 3 and the second equation by -2. c) Multiply the first equation by 4 and the second equation by 3. d) All of the above.

OpenStudy (campbell_st):

decide which variable to eliminate.... x or y...

OpenStudy (campbell_st):

so the question should read to eliminate y, the 1st step should be so if you are wanting to eliminate y... they need the same coefficient with opposite signs I would multiply the first equation by 4 and the 2nd by -3 it would give -8x - 12y = -4 -9x + 12y = 9 then add the equations to eliminate y

OpenStudy (campbell_st):

or if you wanted to eliminate x you would multiply the 1st equation by -3 and the 2nd by 2 or you could even have 1st equation by 3 and the 2nd by -2 so lots of choices.... and it depends on which letter you are going to eliminate 1st

OpenStudy (anonymous):

i honestly have no clue.

OpenStudy (campbell_st):

well a and b will both eliminate x c) is almost correct for eliminating y.... just the that it isn't -4 and 3 or 4 and -3 if its either of those then then answer will be d otherwise a or b

OpenStudy (anonymous):

so if you had to choose one what would it be?

OpenStudy (campbell_st):

I'd choose A